# Learning Robust Anymodal Segmentor with Unimodal and Cross-modal Distillation

### v1

@@ Abstract
Simultaneously using multimodal inputs from multiple sensors to train segmentors is intuitively advantageous but practically challenging. A key challenge is unimodal bias, where multimodal segmentors over-rely on certain modalities, causing performance drops when others are missing—common in real-world applications. To this end, we develop the first framework for learning robust segmentor that can handle any combinations of visual modalities. Specifically, we first introduce a parallel multimodal learning strategy for learning a strong teacher. The cross-modal and unimodal distillation is then achieved in the multiscale representation space by transferring the feature-level knowledge from multimodal to anymodal segmentors, aiming at addressing the unimodal bias and avoiding over-reliance on specific modalities. Moreover, a prediction-level modality-agnostic semantic distillation is proposed to achieve semantic knowledge transferring for segmentation. Extensive experiments on both synthetic and real-world multi-sensor benchmarks demonstrate that our method achieves superior performance (+6.37% & +6.15%).
Figure 1: Comparison between ours against SoTA methods (Zheng et al., 2024b, a) on (a) MUSES and (b) DELIVER datasets.
@@ 1 Introduction
The success of multimodal deep learning relies heavily on effectively leveraging information from multiple modalities, particularly for complex tasks such as semantic segmentation in scene understanding. While intuitively beneficial, training segmentation models with inputs from multiple sensors presents significant practical challenges. A prominent issue in this context is unimodal bias — a phenomenon where networks develop an over-reliance on a single, faster-to-learn modality, often overlooking other sources of valuable information. This bias stems from the distinct characteristics and varied learning dynamics of each sensor modality.
For example, the well-known CMX model (Zhang et al., 2023a) in multimodal semantic segmentation suffers significant performance drops when evaluated without the RGB modality. Meanwhile, the state-of-the-art model Any2Seg (Zheng et al., 2024a), which aims at learning modality-agnostic representation for missing modality problems, demonstrates a significant performance decline when evaluated in modality-incomplete scenarios. Specifically, when depth data is missing, segmentation performance drops markedly (RD: 68.21 $\rightarrow$ R: 39.02, a decrease of 29.19 mIoU), illustrating how unimodal bias can lead to substantial performance degradation in real-world applications where certain modalities are often unavailable.
Despite advancements in multimodal learning, such as leveraging large multimodal language models (Zheng et al., 2024a) and prioritizing each modality (Zheng et al., 2024b), progress in addressing unimodal bias and fostering robust multimodal correlations remains limited. To address this gap, we introduce the first framework for learning robust anymodal segmentors [^1] We define anymodal segmentors as models that ensure robust segmentation performance despite missing modalities.. This framework is tailored to handle real-world scenarios where modality completeness cannot be guaranteed, such as missing modality (Liu et al., 2024) or modality-agnostic segmentation (Zheng et al., 2024a).
Our approach begins with a novel Parallel Multimodal Learning (PML) strategy, which facilitates the learning of a strong teacher model for both unimodal and multimodal distillation without adding extra parameters. Inspired by recent methods (Zheng et al., 2024a, b), we process all multimodal inputs from different sensors in a single mini-batch, passing them through the segmentation model. Modality fusion is performed through simple averaging, and supervision is applied at the final layer of the segmentation decoder. This straightforward yet effective PML strategy enables segmentors to focus on capturing both unimodal and multimodal knowledge, as demonstrated in Tab. 5 and Tab. 6.
@@
Extensive experiments on real-world and synthetic benchmarks demonstrate the superior robustness and performance of our method compared to existing state-of-the-art approaches, achieving mIoU improvements of +6.37% and +6.15%, respectively [^2] https://github.com/zhengxuJosh/AnySeg. Additionally, we analyze why fused multimodal fusion distillation is unsuitable for ensuring robustness in multimodal segmentation.
Figure 2: (a) Overall framework of our proposed method; (b) Unimodal and (c) cross-modal feature distillation.
@@ 2.1 Multimodal Semantic Segmentation
Semantic segmentation with multi-sensor inputs enhances scene understanding by leveraging complementary information from diverse sensors, such as event cameras (Zhou et al., 2024; Zheng & Wang, 2024), LiDAR sensors (Li et al., 2023), and others (Zheng et al., 2023, 2024c, 2024d). Recent advances in multi-sensor systems have led to the development of various approaches (Zheng et al., 2024b, a; Zhang et al., 2023a) and datasets (Zhang et al., 2023b; Brödermann et al., 2024) that extend from dual-modality fusion to full multimodal fusion, with the aim of achieving robust perception across diverse lighting and environmental conditions throughout the day (Broedermann et al., 2023; Wei et al., 2023; Zhang et al., 2021; Man et al., 2023; Wang et al., 2022; Chen et al., 2021; Zhang et al., 2023b, a). For instance, MUSES (Brödermann et al., 2024) dataset integrates data from a frame camera, LiDAR, radar, event camera, and IMU/GNSS sensors to capture driving scenes in adverse conditions with increased uncertainty. Recently, CMNeXt (Zhang et al., 2023b) introduced the task of fusing an arbitrary number of modalities, although this approach still relies primarily on RGB input for optimal performance. In our work, we address the challenge of unimodal bias in multimodal semantic segmentation by focusing on developing a robust anymodal segmentor that can maintain performance across various input combinations, rather than solely optimizing for multimodal segmentation accuracy.
2.2 Missing Modality Robustness
@@
In the multimodal learning community, several studies have sought to understand unimodal bias from both empirical (Kleinman et al., 2023; Peng et al., 2022) and theoretical perspectives (Huang et al., 2022). As shown by Huang et. al (Huang et al., 2022; Zhang et al., 2024), while multimodal learning has the potential to surpass unimodal performance, it often falls short due to modality competition: only the subset of modalities more closely aligned with the encoder’s initial parameters tends to dominate learning within the multimodal network. This phenomenon also occurs in multimodal semantic segmentation, as MAGIC (Zheng et al., 2024b) and Any2Seg (Zheng et al., 2024a) struggle when depth data is missing during inference. However, these studies are limited to relatively simple network architectures, typically using shallow networks, which constrains the applicability of their findings to real-world settings. In this work, we focus on addressing practical challenges in multi-sensor systems that are widely applicable across industrial domains, including autonomous driving and intelligent systems. We define the unimodal bias problems in multimodal semantic segmentation and propose the anymodal framework to evaluate the robustness of multimodal segmentors.
2.3 Robust Multimodal Segmentors
@@
In practical applications, sensor failures or malfunctions often lead to incomplete multimodal data, posing significant challenges for existing segmentation frameworks, which are typically trained and evaluated with full data pairs (Liu et al., 2024). Recent research has focused on developing robust models that, although trained with complete modality sets, can accommodate the random absence of certain modalities during deployment (Liu et al., 2024; Wang et al., 2023b; Maheshwari et al., 2024; Reza et al., 2023; Chen et al., 2023; Zhao et al., 2023). Wang et. al (Wang et al., 2023a) proposed a method to adaptively identify crucial modalities and distill knowledge from them, enabling cross-modal compensation for missing modalities. Liu et. al (Liu et al., 2024) extended this idea by introducing modality-incomplete scene segmentation, addressing both system-level and sensor-level deficiencies. More recently, methods such as MAGIC (Zheng et al., 2024b) and Any2Seg (Zheng et al., 2024a) aim to achieve modality-agnostic segmentation by treating each modality equally and extracting shared, modality-agnostic representations. However, experimental results indicate that the fundamental issue of unimodal bias persists. For example, in the modality-incomplete evaluation of Any2Seg, performance drops significantly when depth data is absent (RD: 68.21 $\rightarrow$ R: 39.02 with a drop of 29.19 mIoU), highlighting the ongoing challenge of unimodal bias in multimodal semantic segmentation. To address this, we propose the first framework for learning robust anymodal segmentors that can withstand real-world scenarios with missing modalities by distilling both multimodal and unimodal knowledge. Additionally, we introduce a parallel multimodal learning strategy as a strong baseline, further advancing robust anymodal segmentation.
@@ 3 Methodology
The overall framework is depicted in Fig. 2. It consists of two segmentors, i.e., the multimodal segmentor $\mathcal{F}_{ms}$ and anymodel segmentor $\mathcal{F}_{as}$, as well as two key modules, including the Unimodal and Multimodal Distillation and the Modality-agnostic Semantic Distillation modules. The mulimodal segmentor $\mathcal{F}_{ms}$ is first pre-trained with our proposed parallel multimodal learning strategy to learn a strong teacher with expertise in multimodal scenarios, its parameter is frozen during training the anymodal segmentor $\mathcal{F}_{as}$.
Inputs. Our framework processes multi-modal visual data from four modalities, all within the same scene. We consider RGB images $\textbf{R}\in\mathbb{R}^{h\times w\times 3}$, depth maps $\textbf{D}\in\mathbb{D}^{h\times w\times C^{D}}$, LiDAR data $\textbf{L}\in\mathbb{L}^{h\times w\times C^{P}}$, and event stack images $\textbf{E}\in\mathbb{E}^{h\times w\times C^{E}}$ to illustrate our method, as depicted in Fig. 2. Here, we follow the data processing as  (Zhang et al., 2023b), where the channel dimensions $C^{D}=C^{P}=C^{E}=3$. Our framework also integrates the corresponding ground truth $Y$ across $K$ categories. For each training iteration, our framework takes a mini-batch $\{r,d,e,l\}$ containing samples from all the input modalities.
@@ 3.1 Parallel Multimodal Learning (PML) Strategy
Recent studies have shown that treating all input modalities equally can enhance multimodal performance in semantic segmentation (Zheng et al., 2024a, b). Building on insights from MAGIC (Zheng et al., 2024b), we adopt a uniform approach for handling all multimodal inputs and introduce a parallel multimodal learning strategy to train a robust teacher model for knowledge distillation. As illustrated in Fig. 3, we compute the mean across multimodal features at each block of the segmentation backbone (Xie et al., 2021). This averaged output serves as the input for the segmentation head, leading to improved multimodal performance, particularly on real-world benchmarks, achieving 51.37 mIoU on MUSES. The supervised loss $L_{pre}$ for training is:
@@
Figure 3: PML for learning a strong multimodal segmentor.
3.2 Unimodal and Multimodal Distillation
@@
Unimodal Distillation. After extracting the multi-scale features $\{f_{r}^{i},f_{d}^{i},f_{e}^{i},f_{l}^{i}\}_{i=1}^{4}$ and $\{g_{r}^{i},g_{d}^{i}\}_{i=1}^{4}$ from the multimodal segmentor (teacher model) and the anymodal segmentor (student model), respectively, we proceed with the distillation process. For the remaining multi-scale features $\{g_{r}^{i},g_{d}^{i}\}_{i=1}^{4}$ from the anymodal segmentor $\mathcal{F}_{am}$, we align them with the corresponding features $\{f_{r}^{i},f_{d}^{i},f_{e}^{i},f_{l}^{i}\}_{i=1}^{4}$ obtained from the multimodal segmentor. The unimodal knowledge distillation loss function based on KL divergency is then formulated as:
$\mathcal{L}_{umd}=\sum_{i=1}^{4}\left(\sum_{j=1}^{C_{i}}g_{r}^{i,j}\log\left(\frac{g_{r}^{i,j}}{f_{r}^{i,j}}\right)+\sum_{j=1}^{C_{i}}g_{d}^{i,j}\log\left(\frac{g_{d}^{i,j}}{f_{d}^{i,j}}\right)\right),$ (4)
where $C_{i}$ denotes the channel count of the $i$-th level features. $\mathcal{L}_{umd}$ facilitates the transfer of knowledge within each modality, thereby enhancing the performance of the anymodal segmentor in handling single-modality data. This is demonstrated in Tab. 5, where we show the performance improvements achieved by applying unimodal distillation in the anymodal segmentation task. However, while unimodal knowledge transfer enhances single-modality performance, especially for the RGB images, it simultaneously hinders multimodal performance when different modality combinations are encountered, as also illustrated in Tab. 5.
Figure 4: Qualitative comparison between CMNeXt (Zhang et al., 2023b), MAGIC (Zheng et al., 2024b) and ours on MUSES.
Table 1: Results of anymodal semantic segmentation validation with three modalities on real-world benchmark MUSES dataset using SegFormer-B0 as backbone model.

### v2

@@ Abstract
Leveraging multimodal inputs from multiple sensors offers intuitive benefits for semantic segmentation but introduces practical challenges—most notably, unimodal bias, where models overfit to dominant modalities and perform poorly when others are missing, a common issue in real-world scenarios. To address this, we propose AnySeg, a unified framework for learning robust segmentors that generalize to arbitrary combinations of input modalities. Our approach first trains a strong multimodal teacher using parallel modality learning. We then distill both unimodal and cross-modal knowledge to an anymodal student via multiscale feature-level distillation, reducing modality dependence and improving generalization. To further enhance semantic consistency, we introduce a prediction-level, modality-agnostic distillation loss. Unlike prior work, our framework explicitly handles missing modalities challenges by learning unimodal and cross-modal correspondence among input modalities. Extensive experiments on synthetic and real-world multi-sensor datasets demonstrate the effectiveness of AnySeg, achieving notable improvements of +6.37% and +6.15% in mIoU.
@@ 1 Introduction
The success of multimodal deep learning relies heavily on effectively leveraging information from multiple modalities, particularly for complex tasks such as semantic segmentation in scene understanding [1, 2, 3, 4]. While intuitively beneficial, training segmentation models, a.k.a., segmentors, with inputs from multiple sensors presents significant practical challenges. A prominent issue in this context is unimodal bias — a phenomenon where networks develop an over-reliance on a single, faster-to-learn modality, often overlooking other sources of valuable information. This bias stems from the distinct characteristics and varied learning dynamics of each sensor modality.
For example, the well-known CMX model [5] in multimodal semantic segmentation suffers significant performance drops when evaluated without the RGB modality. Meanwhile, the state-of-the-art model Any2Seg [6], which aims at learning modality-agnostic representation for missing modality problems, demonstrates a significant performance decline when evaluated in modality-incomplete scenarios. For instance, when depth data is missing, segmentation performance drops markedly (RD: 68.21 $\rightarrow$ R: 39.02, a decrease of 29.19 mIoU), illustrating how unimodal bias can lead to substantial performance degradation in real-world applications where certain modalities are often unavailable.
Despite advancements in multimodal learning, such as leveraging large multimodal language models [6] and prioritizing each modality [7], progress in addressing unimodal bias and fostering robust multimodal correlations remains limited. To address this gap, we introduce the first framework for learning robust anymodal segmentors [^1] We define anymodal segmentors as models that ensure robust performance despite missing modalities.. This framework is tailored to handle real-world scenarios where modality completeness cannot be guaranteed, such as missing modality [8] or modality-agnostic segmentation [6].
Our approach begins with a novel Parallel Multimodal Learning (PML) strategy, which facilitates the learning of a strong teacher model for both unimodal and multimodal distillation without adding extra parameters. Inspired by recent methods [6, 7], we process all multimodal inputs from different sensors in a single mini-batch, passing them through the segmentation backbone, i.e., SegFormer. Multimodal fusion is performed through simple averaging, and supervision is applied at the final layer of the segmentation decoder. This straightforward yet effective PML strategy enables segmentor to focus on capturing both unimodal and multimodal knowledge (See Tab. 5 and Tab. 6).
@@
Extensive experiments on real-world and synthetic benchmarks demonstrate the superior robustness and performance of our method compared to existing state-of-the-art approaches, achieving mIoU improvements of +6.37% and +6.15%, respectively. Moreover, we analyze why fused multimodal fusion distillation is unsuitable for ensuring robustness in multimodal segmentation and further discuss the feature characteristics of multimodal data.
@@ 2 Related Work
Multimodal Semantic Segmentation Semantic segmentation with multi-sensor inputs enhances scene understanding by leveraging complementary information from diverse sensors, such as event cameras [9, 10], LiDAR sensors [11], and others [12, 13, 14, 15, 1]. Recent advances in multi-sensor systems have led to the development of various approaches [16, 7, 6, 5] and datasets [17, 18] that extend from dual-modality fusion to full multimodal fusion, with the aim of achieving robust perception across diverse lighting and environmental conditions throughout the day [19, 20, 21, 22, 23, 24, 25, 17, 5, 26]. For instance, MUSES [18] dataset integrates data from a frame camera, LiDAR, radar, event camera, and IMU/GNSS sensors to capture driving scenes in adverse conditions with increased uncertainty. Recently, CMNeXt [17] introduced the task of fusing an arbitrary number of modalities, although this approach still relies primarily on RGB input for optimal performance. In our work, we address the challenge of unimodal bias in multimodal semantic segmentation by focusing on developing a robust anymodal segmentor that can maintain performance across various input combinations, rather than solely optimizing for multimodal segmentation accuracy.
@@
Missing Modality Robustness In the multimodal learning community, several studies have sought to understand unimodal bias from both empirical [27, 28] and theoretical perspectives [29]. As shown by Huang et. al [29, 30], while multimodal learning has the potential to surpass unimodal performance, it often falls short due to modality competition: only the subset of modalities more closely aligned with the encoder’s initial parameters tends to dominate learning within the multimodal network. This phenomenon also occurs in multimodal semantic segmentation, as MAGIC [7] and Any2Seg [6] struggle when depth data is missing during inference. In this work, we focus on addressing practical challenges in multi-sensor systems that are widely applicable across industrial domains, including autonomous driving and intelligent systems. We define the unimodal bias problems in multimodal semantic segmentation and propose the anymodal semantic segmentation framework.
@@
Robust Multimodal Segmentors. In practice, sensor failures often result in incomplete multimodal data, challenging segmentation frameworks typically trained on complete modality pairs [8]. Recent studies aim to build models that, while trained with full modalities, remain effective when some inputs are missing [8, 31, 32, 33, 34, 35]. Wang et. al [36] proposed adaptive modality selection and knowledge distillation for cross-modal compensation. Liu et. al [8] extended this to modality-incomplete scene segmentation, addressing both system- and sensor-level failures. More recent methods like MAGIC [7] and Any2Seg [6] aim for modality-agnostic segmentation by extracting shared representations. However, unimodal bias remains unresolved. For instance, Any2Seg’s performance drops sharply without depth input (RD: 68.21 $\rightarrow$ R: 39.02), illustrating the challenge. To address this, we propose the first framework for learning robust anymodal segmentors capable of handling missing modalities by distilling both unimodal and cross-modal knowledge. We also introduce a parallel multimodal learning strategy to build a strong teacher model, further advancing robust segmentation under incomplete inputs.
Figure 1: (a) Overall of AnySeg with a two-stage training strategy: the multimodal teacher is first trained using PML, then frozen for student distillation. (b) Unimodal feature distillation transfers intra-modality knowledge. (c) Cross-modal feature distillation enables modality interaction transfer.
@@ 3 Methodology
The overall framework is depicted in Fig. 1. It consists of two segmentors, i.e., the multimodal teacher segmentor $\mathcal{F}_{ms}$ and anymodal student segmentor $\mathcal{F}_{as}$, as well as two key modules, including the unimodal and cross-modal distillation and the modality-agnostic semantic distillation modules. The teacher $\mathcal{F}_{ms}$ is first pre-trained with our proposed parallel multimodal learning strategy to learn a strong teacher with expertise in multimodal scenarios, its parameter is frozen during training the student $\mathcal{F}_{as}$. Inputs: Our framework processes multi-modal visual data from four modalities, all within the same scene. We consider RGB images $\textbf{R}\in\mathbb{R}^{h\times w\times 3}$, depth maps $\textbf{D}\in\mathbb{D}^{h\times w\times C^{D}}$, LiDAR data $\textbf{L}\in\mathbb{L}^{h\times w\times C^{P}}$, and event stack images $\textbf{E}\in\mathbb{E}^{h\times w\times C^{E}}$ to illustrate our method, as depicted in Fig. 1. Here, we follow the data processing as  [17], where the channel dimensions $C^{D}=C^{P}=C^{E}=3$, and we also integrates the corresponding ground truth $Y$ across $K$ categories. For each training iteration, a mini-batch $\{r,d,e,l\}$ contains samples from all the input modalities.
@@ 3.1 Parallel Multimodal Learning (PML) Strategy
Recent studies have shown that treating all input modalities equally can enhance both multimodal and unimodal performance [6, 7]. Building on insights from MAGIC [7], we adopt a uniform approach for handling all multimodal inputs and introduce a parallel multimodal learning strategy to train a robust teacher model for knowledge distillation. As illustrated in Fig. 2, we compute the mean across multimodal features at each block of the segmentation backbone [37]. This averaged output serves as the input for the segmentation head, leading to improved multimodal performance, particularly on real-world benchmarks, achieving 51.37 mIoU on MUSES. The supervised loss $L_{pre}$ for training is:
@@
3.2 Unimodal and Cross-modal Distillation
Figure 2: PML for learning a strong multimodal segmentor as teacher model.
@@ 3.2 Unimodal and Cross-modal Distillation
Unimodal Distillation. After extracting the multi-scale features $\{f_{r}^{i},f_{d}^{i},f_{e}^{i},f_{l}^{i}\}_{i=1}^{4}$ and $\{g_{r}^{i},g_{d}^{i}\}_{i=1}^{4}$ from the multimodal segmentor (teacher model) and the anymodal segmentor (student model), respectively, we proceed with the distillation process. For the remaining multi-scale features $\{g_{r}^{i},g_{d}^{i}\}_{i=1}^{4}$ from the anymodal segmentor $\mathcal{F}_{am}$, we align them with the corresponding features $\{f_{r}^{i},f_{d}^{i},f_{e}^{i},f_{l}^{i}\}_{i=1}^{4}$ obtained from the multimodal segmentor. The unimodal knowledge distillation loss function based on KL divergence is defined as:
$\mathcal{L}_{umd}=\sum_{i=1}^{4}\left(\sum_{j=1}^{C_{i}}\tilde{g}_{r}^{i,j}\log\left(\frac{\tilde{g}_{r}^{i,j}}{\tilde{f}_{r}^{i,j}}\right)+\sum_{j=1}^{C_{i}}\tilde{g}_{d}^{i,j}\log\left(\frac{\tilde{g}_{d}^{i,j}}{\tilde{f}_{d}^{i,j}}\right)\right),$ (4)
where $C_{i}$ denotes the number of channels in the $i$-th level features. To ensure valid probability distributions for KL divergence computation, we apply softmax to the teacher features $g$ and log-softmax to the student features $f$, yielding normalized representations $\tilde{g}$ and $\tilde{f}$. This guarantees non-negativity and unit sum, thus satisfying the theoretical prerequisites of KL divergence. The loss term $\mathcal{L}_{umd}$ promotes intra-modality knowledge transfer, enhancing the anymodal segmentor’s ability to generalize from unimodal inputs. This is demonstrated in Tab. 5, where we show the performance improvements achieved by applying unimodal distillation in the anymodal segmentation task. However, while unimodal knowledge transfer enhances single-modality performance, especially for the RGB images, it simultaneously hinders multimodal performance when different modality combinations are encountered, as also illustrated in Tab. 5.
Cross-modal Correspondence Distillation.
To address this, we leverage cross-modal correspondences between “easy-to-learn” and “hard-to-learn” modalities to achieve a more balanced performance. By distilling these cross-modal relationships from the teacher to the student model—referred to as the “anymodal” segmentor—we effectively mitigate unimodal bias across all modalities. The distillation of cross-modal knowledge between student features $\{g_{r}^{i},g_{d}^{i}\}_{i=1}^{4}$ and teacher features $\{f_{r}^{i},f_{d}^{i},f_{e}^{i},f_{l}^{i}\}_{i=1}^{4}$ is achieved through:
$\mathcal{L}_{cmd}=\sum_{i=1}^{4}\sum_{j=1}^{C_{i}}\tilde{\mathcal{S}}\left(g_{d}^{i,j},g_{r}^{i,j}\right)\log\left(\frac{\tilde{\mathcal{S}}\left(g_{d}^{i,j},g_{r}^{i,j}\right)}{\tilde{\mathcal{S}}\left(f_{d}^{i,j},f_{r}^{i,j}\right)}\right),$ (5)
where $\mathcal{S}(x,y)=\frac{x\cdot y}{\|x\|\|y\|}$ denotes the cosine similarity between feature vectors $x$ and $y$, and $\tilde{\mathcal{S}}(x,y)=\frac{\mathcal{S}(x,y)+1}{2}$ is its normalized form within $[0,1]$ to ensure numerical stability for the logarithmic operation. To further mitigate potential instability caused by negative cosine values—especially when the similarities of teacher and student features differ in sign—we average the similarity scores across batch samples with the same feature shape. Empirically, this averaging leads to non-negative values when modalities are semantically aligned. Nonetheless, we adopt the normalized similarity $\tilde{\mathcal{S}}$ to guarantee theoretical robustness. This formulation aligns cross-modal representations by minimizing the discrepancy between teacher and student similarities across modalities. After applying cross-modal correspondence distillation, the inherent unimodal bias in this task—as well as the bias introduced by unimodal knowledge distillation—are largely mitigated, as shown in Tab. 5.
Figure 3: Qualitative comparison on MUSES.
Table 1: Results of anymodal semantic segmentation validation with three modalities (F: frame camera, E: event cameras, L: LiDAR sensor) on real-world MUSES with SegFormer-B0.